OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 - 2x + 5

OpenStudy (anonymous):

first coordinate of the vertex of \\(y=ax^2+bx+c\) is always \[-\frac{b}{2a}\]

OpenStudy (anonymous):

second coordinate of the vertex is what you get for \(y\) when you replace \(x\) by the first coordinate

OpenStudy (anonymous):

once you have the vertex \((h,k)\) then vertex form is \[y=a(x-h)^2+k\] in your case \(a=1,b=-2\)

OpenStudy (anonymous):

i can check your answer if you like

OpenStudy (anonymous):

y=(x+1)^2-4?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

your signs are messed up

OpenStudy (anonymous):

\[-\frac{b}{2a}=-\frac{-2}{2\times (-1)}=1\]

OpenStudy (anonymous):

if you replace \(x\) by \(1\) you get \(4\) not \(-4\)

OpenStudy (anonymous):

so \[y=(x-1)^2+4\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw